The molarity (M) of an aqueous solution containing 129 g of glucose (C6H12O6) in 200 mL of solution is ________.



1
Expert's answer
2021-10-18T03:09:12-0400

Molarity="\text{moles per litre}molarity" =moles per litre

Number of moles=129/180= 0.717moles

molarity=0.717/200ml= 0.0036M